OpenStudy (thedj4jc):

Solve |2x - 5| = 4. {x | x = -4.5 or x = 4.5} {x | x = 0.5 or x = 4.5} {x | 0.5 < x < 4.5} Please explain how to get answer!!

OpenStudy (anonymous):

Its x = 4.5

OpenStudy (thedj4jc):

Thank you, do you know what all of the other stuff is for. The answer is {x | x = 0.5 or x = 4.5}, but I don't know how they got answer.

OpenStudy (anonymous):

@texaschic101 @perl Im not sure maybe they can help.

OpenStudy (thedj4jc):

Ok, thank you!!

OpenStudy (texaschic101):

| 2x - 5| = 4 2x - 5 = 4 - (2x - 5) = 4 2x = 4 + 5 -2x + 5 = 4 2x = 9 -2x = 4 - 5 x = 9/2 -2x = -1 x = 4 1/2 or 4.5 x = 1/2 or 0.5

OpenStudy (texaschic101):

read over this.. http://www.purplemath.com/modules/solveabs.htm

OpenStudy (thedj4jc):

Thank you!!! That makes so much more sense now!!

OpenStudy (texaschic101):

anytime :)